A body fallung from tha rest has a velocity v after it falls through a height h . The distance it has to fall down further for its velocity to become double will be

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Deepak
 
Use
v^2 – u^2 = 2*g*s
intial velocity u=0
at velocity=v height=h
therefore
u^2 = 2gh
then let intial velocity u=v
and final velocity=2v
therefore
(2v)^2 – v^2 = 2*g*s
4* v^2 – v^2 = 3 v^2 = 6*g*h = 2g* (3h)
the distance it has to fall down further for its velocity to become 2v is 3.times h.
 
Regards
Arun (askIITians forum expert)